OpenStudy (anonymous):

Two times the least of three consecutive odd integers exceeds three times the greatest by 15. How do i set it up to solve?

OpenStudy (anonymous):

if they are consecutive odd integers they differ by 2, you can call them \[x,x+2,x+4\]

OpenStudy (anonymous):

then two times the first one is \(2x\) which exceeds 3 times the last one \(3(x+4)\) by \(16\)

OpenStudy (anonymous):

oops i meant by 15 not 16

OpenStudy (anonymous):

that means \[2x-15=3(x+4)\] or you could write \[2x=3(x+4)+15\]

OpenStudy (anonymous):

lets go with the first one \[2x-15=3(x+4)\]

OpenStudy (anonymous):

Ok thanks a million! @satellite73

OpenStudy (anonymous):

multiply out and get \[2x-15=3x+12\] then solve for \(x\)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so x=-27 @satellite73

OpenStudy (anonymous):

that is what i get, yes we can check it if you like

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

i got -69=-69

OpenStudy (anonymous):

they will be \(-27,-25,-23\) and \[2\times -27=-54\] \[3\times -23=-69\] and \(-54\) exceeds \(-69\) by \(15\) since \(-69+15=-54\)

OpenStudy (anonymous):

looks like it is right

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw